• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Wendy

    Log entry time 16:19:16 on December 3,2001

    Entry number 78685

    keyword=Shift Summary

    8:50 Mcc took the beam to tune
    9:00 adaql4 has frozen, Bob Michaels will come in later and reboot the machine
    9:40 Mcc will leave the OTR 1C12 in to continue to monitor the beam's stablity
    9:50 Beam has returned
    MCC is still looking into the beam instablity but they are delivering beam at 76 uA until they are able to locate the problem
    10:45 Adaql4 has been rebooted.
    1:30 Beam gone
    2:00 Beam is back
    2:30 beam gone , Mcc will cycle a box supply ~10 min.
    Shift was slow. There are problems with the beam
    Total 1087.44 uA.hour for J+